OpenStudy (anonymous):

-18+4+x=9-19? Will medal and fan.

OpenStudy (foreverandalways13):

do you need to solve for x?

OpenStudy (anonymous):

Yes.

OpenStudy (ccastudentgrade10):

x = 4 i believe

OpenStudy (anonymous):

4 i think

OpenStudy (anonymous):

Explain please.

OpenStudy (foreverandalways13):

hey guys help him through it not just give him the answer

OpenStudy (anonymous):

Thank you! @foreverandalways13

OpenStudy (foreverandalways13):

oh no problem

OpenStudy (anonymous):

-18+4 =14 and 9-19= -10 and 14 - 10 = 4 :D

OpenStudy (anonymous):

Thanks! Can you guys help on a few more?

OpenStudy (anonymous):

Thank you for the medals!!! :D

OpenStudy (foreverandalways13):

sure

OpenStudy (anonymous):

maybe idk what grade this is I am in 6th

OpenStudy (anonymous):

9th

OpenStudy (anonymous):

And 10th. I'm in 8th, but I go ahead of my curriculum to make sure I understand things later on.

OpenStudy (anonymous):

o well,i might be able to help still.I will try,but if i am not sure i will say "I think" okay? :D

OpenStudy (foreverandalways13):

12th

OpenStudy (foreverandalways13):

i can help!!

OpenStudy (anonymous):

Thanks, new question to prevent spam and ban.

OpenStudy (foreverandalways13):

ok

OpenStudy (ccastudentgrade10):

if you want i can help too